Vice President Dickson,

The recently announced plan to remove the Fishbowl Athena Cluster from its current location should be immediately reconsidered. Students have exhibited great support for the Fishbowl, and we believe moving it would be a serious error. We also believe the current location better serves students than the planned alternatives. We urge the President, the Planning Office and the Student Services Reengineering Team to maintain the Fishbowl in its present location.

Yours truly,

Matthew Gray, Yonah Schmeidler, and the body of the Student Information Processing Board